OpenStudy (anonymous):

Solve the following system of equations: x - 2y = 14 x + 3y = 9 A) (1, 12) B) (-1, -12) C) (12, -1) D) (12, 1)

OpenStudy (anonymous):

Remember what I taught you about plugging in for x and y with each answer @jut21?

OpenStudy (anonymous):

not really...

OpenStudy (anonymous):

This was your first problem: Solve the following system of equations: -2x + y = 1 -4x + y = -1 A) (3, 1) B) (-1, 3) C) (-1, -3) D) (1, 3) And this is what I showed you to do: A -2(3) + 1 = 1 -6 + 1 = 1 -5 = 1 (That is not right CROSS A OFF) B -2(-1) + 3 = 1 2 + 3 = 1 5 = 1 (That is not right CROSS B OFF) C -2(-1) + (-3) = 1 2 - 3 = 1 -1 = 1 (That is not right CROSS C OFF) D -2(1) + 3 = 1 -2 + 3 = 1 1 = 1 (Yes, move on to the next) -4(1) + (3) = -1 -4 + 3 = -1 -1 = -1 YOUR ANSWER IS D. Now can you try to do what I did. Plugging in with answers,

OpenStudy (anonymous):

can you do it for me? your really good at these

OpenStudy (anonymous):

No, I am sorry you have to participate too. No offense, people can't just do the work for you. Best way to learn is to physically do it. Now it is really easy almost like elementary school stuff just plug in the x and y's for each answer.

OpenStudy (anonymous):

So what did you get for A @jut21?